OpenStudy (anonymous):

solve the following equation for A: 2A/3=8+4A

OpenStudy (amistre64):

get all the A stuff to one side... (2a/3) -4a = 8

OpenStudy (amistre64):

get like denominators on the left side to make them add-able

OpenStudy (anonymous):

ok im doing this without a calculator i cnt use one because im going in the military and its an asvab test everything i do has to be on paper

OpenStudy (amistre64):

(2a/3) - (12a/3) = 8 -10a / 3 = 8

OpenStudy (amistre64):

i hardly ever use a calculator :)

OpenStudy (anonymous):

me and math not a good combination

OpenStudy (amistre64):

we can multiply this by (3) to get a on solid ground: (3)(-10a/3) = 8(3) -10a = 24 .... you with me so far?

OpenStudy (anonymous):

um kinda slow down just a little bit

OpenStudy (amistre64):

let me know if anything I have done doesnt make any sense to you

OpenStudy (anonymous):

at the top (2a/3)-12a/3=8 how did you get 12 did u multiply the 3 and 4

OpenStudy (amistre64):

yes. 4a (3) 12a --- x --- = ---- 1 (3) 3

OpenStudy (anonymous):

so so far im with you

OpenStudy (amistre64):

we should be at: -10a = 24 to get that extra baggage (-10) off of the "a"; we divide it by (-10), becasue -10/-10 = 1. But what we do to one side HAS to be done to the other to keep them EQUAL. (-10a)/-10 = 24/-10 a = -24/10. How good are you at reducing?

OpenStudy (anonymous):

one moment let me write this down

OpenStudy (anonymous):

ok (-10a)/-10=24 how

OpenStudy (amistre64):

You miss typed it; it is: (-10a)/-10 = 24/-10 -10a 24 ---- = --- since -10/-10 = 1 we can write this this way: -10 -10 24 (1)a = --- -10

OpenStudy (anonymous):

Ok one moment

OpenStudy (anonymous):

ok now what do we do

OpenStudy (amistre64):

now we have "a" all by its little lonesome and can use it as it is; OR as some tests do, we can "reduce" the fraction, and or put it into decinal form. Its good to know how to do any of them.

OpenStudy (amistre64):

what I do to reduce; is I see if the numbers are even and simply halve them till I cant halve no more...

OpenStudy (anonymous):

example please

OpenStudy (amistre64):

-24/2 -12 ------ = ---- 10/2 5 -12/5 is as far as itll go and there and 5 cannot divide into 12 evenly so we can leave it like this: -12/5

OpenStudy (anonymous):

ya know i feel really dumb the whole time i was thinking that this symbol / meant to divide wow blonde moment

OpenStudy (amistre64):

it is a fraction bar, and all fractions are are divisions of whole numbers.... so yes, it does mean divide :)

OpenStudy (anonymous):

oh ok good i thought i was loosing my mind

OpenStudy (amistre64):

just be able to see it for all of its potential :)

OpenStudy (anonymous):

ok we can keep going :D i want to thank you for helping me. i have a learning disability and math without a calculator is like finding a needle in a hay stack

OpenStudy (amistre64):

me.... im just senile :)

OpenStudy (anonymous):

hahahaha do you have facebook so if i have any more ? i can ask you

OpenStudy (anonymous):

lets keep going

OpenStudy (amistre64):

to keep going, we found the value for "a" already: a = -12/5 If we want to express this as a "mixed" number; we simply do the division and set the "remainder" back over (5) like this: 2 <- we have a whole number of (2) ----- 5| 12 -10 --- 2 <-- we have a remainder of (2), so put it back on the (5) our mixed number for "a" is: a = 2 2/5

OpenStudy (anonymous):

ok hold on while i write it down :D

OpenStudy (amistre64):

i forgot the (-) sign... oops make that: a = -2 2/5

OpenStudy (anonymous):

ok im back

OpenStudy (amistre64):

do you understand it all so far?

OpenStudy (anonymous):

kinda

OpenStudy (amistre64):

the last thing I can try to teach you is how to turn this into a decimal. If we want to turn -12/5 into a decimal, than we simply keep dividing it til we run out of numbers to have as a remainder: For example: 2.4 -------- 5| 12.000 <- add some "0"s to use. -10 --- 20 <- bring one of the "0"'s down to continue our division -20 ---- 0 <-- nothing left to play with, its a "0". our "decimal" form for the value of "a" is: a = -2.4 Now the library closes soon, so I have to go home. Ciao :)

OpenStudy (anonymous):

ok thank you so much for your help